7: find wavenumber, wavelength and frequency of Ha line of Hydrogen, assuming that the nucleus has infinite mass. Calculate the wavelength of the Balmer series limit.


1
Expert's answer
2021-02-16T10:11:58-0500

Wave number is reciprocal of wavelength i.e. number of wavelength per centimeter. It is denoted by "\bar{v}". H-alpha (Hα​) is a specific deep-red visible spectral line in the Balmer series; it occurs when a hydrogen electron falls from its third to second lowest energy level. For Hα​ - line in Balmer series of hydrogen spectrum "\bar{v}" is given by "\bar{v}" = R["\frac{1}{2^2}" - "\frac{1}{3^2}" ] = "\frac{5R}{36}" = 1.52 "\times" 106, where R is Rydberg constant.


Wavelength ="\frac{1}{wawe №}" = 656.28 nm.


Frequency:

ΔE = hf = −13.6("\frac{1}{9} - \frac{1}{4}") eV

ΔE = (13.6 "\times" 5 / 36) "\times" 1.6 "\times" 10−19 J

f = "\frac{ΔE}{h}" ​= "\frac{13.6 \times 5 \times 1.6 \times 10−19}{36 \times 6.6 \times 10−34}" = 4.57 "\times" 1014 Hz


For series limit of Balmer series,

p = 2 and n = ∞

"\frac{}{}""\frac{1}{ λ}" = R("\frac{1}{p^2}-\frac{1}{n^2}") = R("\frac{1}{4} - \frac{1}{∞}")

∴λ = "\frac{4}{R}"